# Dynamical Geometric Theory of Principal Bundle Constrained Systems: Strong Transversality Conditions and Variational Framework for Gauge Field Coupling

### v1

@@ Abstract
This paper establishes a complete geometric mechanics framework for constrained systems on principal bundles. The core contributions are the proposal and rigorous characterization of the strong transversality condition, and the dynamic connection equations derived from variational principles. The strong transversality condition establishes an intrinsic connection between constraint distributions and principal bundle connection structures by introducing a Lie algebra dual distribution function $\lambda:P\to\mathfrak{g}^{*}$. We rigorously prove that this is equivalent to a $G$-equivariant splitting of the Atiyah exact sequence. Using differential geometry methods, we prove the existence theorem for the $\lambda$ function (for principal bundles satisfying $\text{ad}^{*}_{\Omega}\lambda=0$ with parallelizable base manifolds) and uniqueness theorem (for semi-simple Lie algebras with trivial centers), providing a theoretical foundation for the geometric classification of constrained systems. From variational principles, we derive the dynamic connection equation $\partial_{t}\omega=d^{\omega}\eta-\iota_{X_{H}}\Omega$, which reveals the interaction mechanism between gauge fields and matter fields in constrained systems, and maintains gauge covariance supported by the decomposition theorem of symplectic potentials and connections. Additionally, we construct the Spencer cohomology mapping and prove its isomorphism relationship with the de Rham cohomology of the principal bundle, establishing an exact correspondence between topological invariants of constrained systems and physical conservation laws. Through the application of Zorn’s lemma, we provide a constructive proof of hierarchical fibrization, explaining the topological mechanism of constraint structure changes. In terms of theoretical applications, we establish a principal bundle formulation for two-dimensional ideal fluids, analyze the geometric correspondence between Spencer characteristic classes and Kelvin’s circulation theorem, and discuss the connection between the constrained Hamiltonian $H=\int_{M}\text{tr}(\omega\wedge\star\Omega)$ and Yang-Mills theory. We also compare the performance differences between standard transversality and strong transversality conditions in non-holonomic constrained systems, revealing that the strong transversality condition can capture geometric effects caused by constraint-curvature coupling. The theoretical framework of this paper unifies gauge field theory and constrained mechanics, providing a rigorous mathematical foundation for understanding non-ideal constraints, geometric phases, and topological invariants. Future work will include extending the theory to infinite-dimensional cases, studying quantization, developing characteristic class-preserving algorithms, and experimental verification in fluid dynamics and non-holonomic systems, which will provide new theoretical tools for analyzing complex physical systems.
@@ 1 Introduction
Constrained systems on principal bundles constitute a fundamental mathematical structure, providing a unified geometric framework for gauge field theories, mechanical systems, and geometric control theory. This structure not only reveals the intrinsic symmetry of physical laws from the perspective of Lie group actions but also provides precise mathematical representations for various physical phenomena—from Yang-Mills field theory to non-holonomic mechanical systems. This paper is dedicated to developing a rigorous mathematical framework for describing and analyzing constrained systems on principal bundles, with a particular focus on the differential geometric characterization of strong transversality conditions and the resulting dynamic connection equations.
@@
The fundamental importance of the differential geometric framework in the study of constrained systems is reflected in the connection theory of Kobayashi and Nomizu[8] and the exact sequence construction of Atiyah[9]. The latter provides an algebraic description for understanding the differential structure of principal bundles and has profound applications in gauge field theory. Kolar et al.[10] and Montgomery[11] examined the application of connection theory in non-holonomic mechanics, but mainly focused on the standard transversality condition—where the constraint distribution and the fiber direction form a direct sum. In recent years, Blohmann[12] studied constrained Hamiltonian systems on principal bundles, Fernandes[13] systematically analyzed the geometric properties of non-integrable distributions, and Cortés et al.[14] developed a non-holonomic mechanics framework within Lie groupoids. However, these works have not systematically explored the dynamic coupling relationship between constraint distributions and principal bundle curvature.
@@
The theoretical innovation of this paper lies in introducing and rigorously characterizing the strong transversality condition, which is not only stronger than the traditional standard transversality condition but also establishes an intrinsic connection with the principal bundle connection form through the Lie algebra dual distribution function $\lambda:P\rightarrow\mathfrak{g}^{*}$. Unlike the momentum map in the Marsden-Ratiu framework or Bloch’s non-holonomic theory, our strong transversality condition requires the distribution function $\lambda$ to satisfy the modified Cartan equation $d\lambda+\text{ad}^{*}_{\omega}\lambda=0$, making it a covariant constant under the connection $\omega$. This goes beyond the standard principal bundle constraint theory of Kolar et al., which only considers the horizontal-vertical decomposition of connections without emphasizing covariant parallelism constraints. The strong transversality condition is mathematically equivalent to a specific $G$-equivariant splitting of the Atiyah exact sequence, and physically corresponds to the dynamic coupling between gauge fields and constrained systems. This coupling has fundamental significance in quantum chromodynamics, superconductor theory, and nonlinear fluid systems, with its essence being the geometric interaction between the curvature form $\Omega$ and the Lie algebra dual distribution $\lambda$.
@@
From a physical perspective, the strong transversality condition and the modified Cartan equation $d\lambda+\text{ad}^{*}_{\omega}\lambda=0$ together describe how constraint forces couple with physical systems through the curvature of gauge fields. While Blohmann’s work[12] examined the Hamiltonian structure of constrained systems on principal bundles, it did not systematically explore the explicit relationship between constraint forces and curvature. Our framework reveals why certain constrained systems (such as magnetofluids) exhibit dynamical behavior different from purely geometric constraints—constraint forces not only restrict motion but also actively participate in energy and momentum exchange, a phenomenon not fully explained in Fernandes’ integrability analysis[13] or traditional non-holonomic theory. In the context of Yang-Mills theory, this corresponds to the non-trivial interaction between gauge fields and matter fields; in fluid dynamics, it describes how the topological properties of vortices influence overall flow patterns.
@@
The rigorous mathematical treatment in this paper covers several key aspects: First, we establish a complete analysis of the Lie algebra dual distribution function $\lambda$ under the strong transversality condition framework, rigorously proving existence and uniqueness theorems under conditions of trivial center ($\mathfrak{z}(\mathfrak{g})=0$) and parallelizable base manifold; Second, we rigorously derive the dynamic connection equation $\partial_{t}\omega=d^{\omega}\eta-\iota_{X_{H}}\Omega$ from variational principles, and prove the exact relationship between the geometric representation of constraint forces and gauge field curvature $P_{\text{constraint}}=\langle\lambda,\Omega(\dot{q},\dot{q})\rangle$—a relationship not explicitly revealed in Marsden and Ratiu’s constrained Hamiltonian theory[7] or Cortés et al.’s work[14]; Third, we rigorously address spectral sequence convergence issues, proving that $d_{2}=0$ under the conditions of compact parallelizable base manifold and compact semi-simple structure group, establishing the relationship between Spencer cohomology and de Rham cohomology, and precisely characterizing topological obstacles in the non-flat principal bundle case; Finally, through the application of Zorn’s lemma, we rigorously construct hierarchical fibrization structures, proving the existence of a global decomposition $P=\bigcup_{\alpha}P_{\alpha}$, achieving an exact topological classification of constrained state spaces.
@@ 3.1.2 Basic Definition and Intuitive Explanation
The strong transversality condition establishes a precise relationship between constraints and connection structures by introducing a Lie algebra dual distribution function:
@@
Given a constraint distribution $\mathcal{D}\subset TP$ on a principal bundle $P\stackrel{{\scriptstyle\pi}}{{\to}}M$, we say $\mathcal{D}$ satisfies the strong transversality condition if there exists a smooth function $\lambda:P\to\mathfrak{g}^{*}$ such that:
@@
    $\mathcal{D}_{p}=\{v\in T_{p}P\ |\ \langle\lambda(p),\omega(v)\rangle=0\}$
@@
    $d\lambda+\mathrm{ad}^{*}_{\omega}\lambda=0$ (modified Cartan structure equation)
where $\omega$ is the connection form, and $\lambda$ is the Lie algebra dual distribution function.
Intuitively, the first part of the strong transversality condition requires that the constraint distribution $\mathcal{D}$ can be represented as a ”matching zero set” of the connection $\omega$ and distribution function $\lambda$. At each point $p$, the constraint plane $\mathcal{D}_{p}$ is defined by the hyperplane satisfying $\langle\lambda(p),\omega(v)\rangle=0$, where $\lambda(p)$ determines the ”orientation” or ”normal vector” of the hyperplane. Compared with the standard transversality condition, this provides more precise control over the geometric shape of the constraint distribution.
The second condition $d\lambda+\mathrm{ad}^{*}_{\omega}\lambda=0$ requires that the distribution function $\lambda$ and the connection $\omega$ satisfy a special covariant parallel relationship. This ensures geometric consistency of the constraint distribution during system evolution, especially when the system moves along the curvature direction of the gauge field.
@@ 3.2.1 Geometric Meaning of the Modified Cartan Equation
The modified Cartan equation in the strong transversality condition:
@@
    This indicates that $\lambda$ is a covariant constant relative to the connection $\omega$, i.e., for any horizontal vector field $X\in\Gamma(\mathcal{H})$:
@@
    Geometrically, this means that when translated along the horizontal direction of the connection, the change in $\lambda$ is exactly compensated by the adjoint action, ensuring that the constraint distribution $\mathcal{D}$ remains invariant. This can be analogized to covariant constant tensor fields under parallel transport, but specifically adapted to the geometric structure of constrained systems.
@@
    This ensures that the constraint distribution $\mathcal{D}$ remains invariant under gauge transformations, reflecting the mechanism of maintaining gauge symmetry in physical systems.
@@
    Curvature Relations and Integrability Conditions: Let the curvature form of the principal bundle be $\Omega=d\omega+\frac{1}{2}[\omega,\omega]$. Applying exterior derivative to the modified Cartan equation:
@@
    $0$ $=\mathrm{ad}^{*}_{\Omega}\lambda\quad(\text{when}\,d\lambda+\mathrm{ad}^{*}_{\omega}\lambda=0)$
    This leads to the key condition $\mathrm{ad}^{*}_{\Omega}\lambda=0$, indicating that the coadjoint action of curvature $\Omega$ on $\lambda$ vanishes. This condition is directly related to the Frobenius integrability of the constraint distribution $\mathcal{D}$. Specifically, for any $X,Y\in\Gamma(\mathcal{D})$:
@@
    $=\langle\lambda,\Omega(X,Y)\rangle\quad(\text{because}\,\omega(X)=\omega(Y)=0\,\text{in}\,\mathcal{D}\cap\mathcal{H})$
    $=0\quad(\text{from the condition}\,\mathrm{ad}^{*}_{\Omega}\lambda=0)$
    Therefore $X,Y]\in\Gamma(\mathcal{D})$, indicating that the constraint distribution forms an integrable distribution under specific conditions.
@@
    Momentum Mapping Interpretation: From a symplectic geometry perspective, $\lambda:P\to\mathfrak{g}^{*}$ can be understood as a momentum mapping, satisfying:
    $d\langle\lambda,\xi\rangle=\iota_{\xi_{P}}\omega,\quad\forall\xi\in\mathfrak{g}$
    where $\xi_{P}$ is the fundamental vector field generated by $\xi$. The modified Cartan equation corresponds to the equivariance condition of the momentum mapping, revealing the deep compatibility between $\lambda$ and the $G$-action on the principal bundle. In physical systems, this corresponds to the balance between constraint forces and gauge fields, manifesting as the geometric foundation of the ”principle of no response to constraints”.
Remark 3.
The modified Cartan equation has profound physical significance: it ensures that the constraint distribution $\mathcal{D}$ maintains structural consistency during system evolution, keeping constraint conditions dynamically coordinated with the geometric characteristics of the system. This mechanism explains why constrained systems under strong transversality conditions can generate geometrically self-consistent dynamical behavior, especially in cases involving gauge fields.
3.2.2 Equivalence of Strong Transversality Condition and Atiyah Exact Sequence Splitting
To understand the deep geometric meaning of the strong transversality condition, we first need to introduce the concept of fundamental vector fields related to principal bundle structures:

### v2

@@ Abstract
This paper establishes a complete geometric mechanics framework for constrained systems on principal bundles. The core contributions are the proposal and rigorous characterization of the strong transversality condition through compatible pairs, and the dynamic connection equations derived from variational principles. The strong transversality condition establishes an intrinsic connection between constraint distributions and principal bundle connection structures by introducing compatible pairs $(\mathcal{D},\lambda)$ where $\lambda:P\to\mathfrak{g}^{*}$ is a Lie algebra dual distribution function. We rigorously prove that this is equivalent to a $G$-equivariant splitting of the Atiyah exact sequence. Using differential geometry methods, we prove the existence theorem for compatible pairs (for principal bundles satisfying $\text{ad}^{*}_{\Omega}\lambda=0$ with parallelizable base manifolds) and uniqueness theorem (for semi-simple Lie algebras with trivial centers), providing a theoretical foundation for the geometric classification of constrained systems. From variational principles, we derive the dynamic connection equation $\partial_{t}\omega=d^{\omega}\eta-\iota_{X_{H}}\Omega$, which reveals the interaction mechanism between gauge fields and matter fields in constrained systems, and maintains gauge covariance supported by the decomposition theorem of symplectic potentials and connections. Additionally, we construct the Spencer cohomology mapping and prove its isomorphism relationship with the de Rham cohomology of the principal bundle, establishing an exact correspondence between topological invariants of constrained systems and physical conservation laws. Through the application of Zorn’s lemma, we provide a constructive proof of hierarchical fibrization, explaining the topological mechanism of constraint structure changes. In terms of theoretical applications, we establish a principal bundle formulation for two-dimensional ideal fluids, analyze the geometric correspondence between Spencer characteristic classes and Kelvin’s circulation theorem, and discuss the connection between the constrained Hamiltonian $H=\int_{M}\text{tr}(\omega\wedge\star\Omega)$ and Yang-Mills theory. We also compare the performance differences between standard transversality and strong transversality conditions in non-holonomic constrained systems, revealing that the strong transversality condition can capture geometric effects caused by constraint-curvature coupling. The theoretical framework of this paper unifies gauge field theory and constrained mechanics, providing a rigorous mathematical foundation for understanding non-ideal constraints, geometric phases, and topological invariants. Future work will include extending the theory to infinite-dimensional cases, studying quantization, developing characteristic class-preserving algorithms, and experimental verification in fluid dynamics and non-holonomic systems, which will provide new theoretical tools for analyzing complex physical systems.
@@ 1 Introduction
Constrained systems on principal bundles constitute a fundamental mathematical structure, providing a unified geometric framework for gauge field theories, mechanical systems, and geometric control theory. This structure not only reveals the intrinsic symmetry of physical laws from the perspective of Lie group actions but also provides precise mathematical representations for various physical phenomena–from Yang-Mills field theory to non-holonomic mechanical systems. This paper is dedicated to developing a rigorous mathematical framework for describing and analyzing constrained systems on principal bundles, with a particular focus on the differential geometric characterization of strong transversality conditions through compatible pairs and the resulting dynamic connection equations.
@@
The fundamental importance of the differential geometric framework in the study of constrained systems is reflected in the connection theory of Kobayashi and Nomizu[8] and the exact sequence construction of Atiyah[9]. The latter provides an algebraic description for understanding the differential structure of principal bundles and has profound applications in gauge field theory. Kolar et al.[10] and Montgomery[11] examined the application of connection theory in non-holonomic mechanics, but mainly focused on the standard transversality condition–where the constraint distribution and the fiber direction form a direct sum. In recent years, Blohmann[12] studied constrained Hamiltonian systems on principal bundles, Fernandes[13] systematically analyzed the geometric properties of non-integrable distributions, and Cortés et al.[14] developed a non-holonomic mechanics framework within Lie groupoids. However, these works have not systematically explored the dynamic coupling relationship between constraint distributions and principal bundle curvature.
@@
The theoretical innovation of this paper lies in introducing and rigorously characterizing the strong transversality condition through the framework of compatible pairs, which is not only stronger than the traditional standard transversality condition but also establishes an intrinsic connection with the principal bundle connection form through compatible pairs $(\mathcal{D},\lambda)$ where $\lambda:P\rightarrow\mathfrak{g}^{*}$ is the Lie algebra dual distribution function. Unlike the momentum map in the Marsden-Ratiu framework or Bloch’s non-holonomic theory, our strong transversality condition requires that $(\mathcal{D},\lambda)$ form a compatible pair satisfying both the compatibility condition $\mathcal{D}_{p}=\{v\in T_{p}P:\langle\lambda(p),\omega(v)\rangle=0\}$ and the modified Cartan equation $d\lambda+\text{ad}^{*}_{\omega}\lambda=0$, making $\lambda$ a covariant constant under the connection $\omega$. This goes beyond the standard principal bundle constraint theory of Kolar et al., which only considers the horizontal-vertical decomposition of connections without emphasizing covariant parallelism constraints. The strong transversality condition is mathematically equivalent to a specific $G$-equivariant splitting of the Atiyah exact sequence, and physically corresponds to the dynamic coupling between gauge fields and constrained systems. This coupling has fundamental significance in quantum chromodynamics, superconductor theory, and nonlinear fluid systems, with its essence being the geometric interaction between the curvature form $\Omega$ and the Lie algebra dual distribution $\lambda$.
@@
From a physical perspective, the compatible pair structure and the modified Cartan equation $d\lambda+\text{ad}^{*}_{\omega}\lambda=0$ together describe how constraint forces couple with physical systems through the curvature of gauge fields. While Blohmann’s work[12] examined the Hamiltonian structure of constrained systems on principal bundles, it did not systematically explore the explicit relationship between constraint forces and curvature. Our compatible pair framework reveals why certain constrained systems (such as magnetofluids) exhibit dynamical behavior different from purely geometric constraints–constraint forces not only restrict motion but also actively participate in energy and momentum exchange, a phenomenon not fully explained in Fernandes’ integrability analysis[13] or traditional non-holonomic theory. In the context of Yang-Mills theory, this corresponds to the non-trivial interaction between gauge fields and matter fields; in fluid dynamics, it describes how the topological properties of vortices influence overall flow patterns.
@@
The rigorous mathematical treatment in this paper covers several key aspects: First, we establish a complete analysis of compatible pairs under the strong transversality condition framework, rigorously proving existence and uniqueness theorems under conditions of trivial center ($\mathfrak{z}(\mathfrak{g})=0$) and parallelizable base manifold; Second, we rigorously derive the dynamic connection equation $\partial_{t}\omega=d^{\omega}\eta-\iota_{X_{H}}\Omega$ from variational principles applied to compatible pairs, and prove the exact relationship between the geometric representation of constraint forces and gauge field curvature $P_{\text{constraint}}=\langle\lambda,\Omega(\dot{q},\dot{q})\rangle$–a relationship not explicitly revealed in Marsden and Ratiu’s constrained Hamiltonian theory[7] or Cortés et al.’s work[14]; Third, we rigorously address spectral sequence convergence issues, proving that $d_{2}=0$ under the conditions of compact parallelizable base manifold and compact semi-simple structure group, establishing the relationship between Spencer cohomology and de Rham cohomology, and precisely characterizing topological obstacles in the non-flat principal bundle case; Finally, through the application of Zorn’s lemma, we rigorously construct hierarchical fibrization structures, proving the existence of a global decomposition $P=\bigcup_{\alpha}P_{\alpha}$, achieving an exact topological classification of constrained state spaces.
@@ 3.1.2 Basic Definition and Intuitive Explanation
The strong transversality condition establishes a precise relationship between constraints and connection structures through the concept of compatible pairs, avoiding circular definitions.
Definition 14 (Compatible Pair).
Given a principal bundle $P\stackrel{{\scriptstyle\pi}}{{\to}}M$ with connection form $\omega$, a pair $(\mathcal{D},\lambda)$ is called a compatible pair if:
  1. 1.
    $\mathcal{D}\subset TP$ is a smooth distribution
  2. 2.
    $\lambda:P\to\mathfrak{g}^{*}$ is a smooth function
  3. 3.
    Compatibility condition: $\mathcal{D}_{p}=\{v\in T_{p}P\mid\langle\lambda(p),\omega(v)\rangle=0\}$
  4. 4.
    Differential condition: $d\lambda+\mathrm{ad}^{*}_{\omega}\lambda=0$ (modified Cartan structure equation)
@@
A constraint distribution $\mathcal{D}\subset TP$ satisfies the strong transversality condition if there exists a smooth function $\lambda:P\to\mathfrak{g}^{*}$ such that $(\mathcal{D},\lambda)$ forms a compatible pair.
Equivalently, we say that $\lambda:P\to\mathfrak{g}^{*}$ is a strong transversal map if there exists a distribution $\mathcal{D}$ such that $(\mathcal{D},\lambda)$ forms a compatible pair.
Theorem 4 (Forward Construction: From $\lambda$ to $\mathcal{D}$).
Let $\lambda:P\to\mathfrak{g}^{*}$ satisfy the modified Cartan equation $d\lambda+\mathrm{ad}^{*}_{\omega}\lambda=0$. Define $\mathcal{D}_{p}=\{v\in T_{p}P\mid\langle\lambda(p),\omega(v)\rangle=0\}$.
Then $(\mathcal{D},\lambda)$ forms a compatible pair, and $\mathcal{D}$ satisfies the strong transversality condition.
Theorem 5 (Inverse Construction: From $\mathcal{D}$ to $\lambda$ (Variational Method)).
Let $\mathcal{D}\subset TP$ be a given constraint distribution satisfying:
@@
    $\mathcal{D}$ is $G$-invariant: $R_{g*}(\mathcal{D}_{p})=\mathcal{D}_{pg}$
@@
    Transversality: $\mathcal{D}_{p}\cap V_{p}=\{0\}$ (where $V_{p}$ is the vertical subspace)
  3. 3.
    Constant rank: $\dim\mathcal{D}_{p}=\dim M+\dim G-k$ for some constant $k$
Define the compatibility functional:
$I[\lambda]=\frac{1}{2}\int_{P}|d\lambda+\mathrm{ad}^{*}_{\omega}\lambda|^{2}+\alpha\int_{P}\mathrm{dist}^{2}(\lambda(p),\mathcal{D}_{p}^{\perp})$
Then there exists $\lambda^{*}$ minimizing $I[\lambda]$ such that $(\mathcal{D},\lambda^{*})$ forms a compatible pair. Therefore $\mathcal{D}$ satisfies the strong transversality condition.
Theorem 6 (Equivalence Theorem).
Let $(\mathcal{D},\lambda)$ be a compatible pair. Then:
  1. 1.
    $\lambda$ is a minimizer of the functional $I[\lambda]$ in Theorem 5
  2. 2.
    $\mathcal{D}$ is the distribution determined by $\lambda$ in Theorem 4
Geometric Intuition: The compatible pair framework provides two complementary perspectives:
  • Constraint-driven approach: Given physical constraints $\mathcal{D}$, find the optimal dual function $\lambda$ that encodes the constraint geometry
  • Dual-driven approach: Given a dual function $\lambda$ satisfying differential compatibility, construct the corresponding constraint distribution $\mathcal{D}$
The strong transversality condition ensures that these two approaches are equivalent and yield the same geometric structure.
Remark 2 (Existence and Well-definedness).
The existence of compatible pairs depends on both the topological properties of the base manifold and the differential geometric structure of the principal bundle. The variational approach in Theorem 5 provides constructive existence criteria under appropriate geometric assumptions, validating the well-definedness of this framework.
Remark 3 (Geometric Motivation).
The modified Cartan equation $d\lambda+\mathrm{ad}^{*}_{\omega}\lambda=0$ can be interpreted as a generalized parallel transport condition: it ensures that the constraint structure is preserved under the natural geometric flows defined by the connection. This leads to a rich interplay between constraint geometry and gauge field curvature, which forms the foundation for the subsequent geometric analysis.
@@ 3.2.1 Geometric Meaning of the Modified Cartan Equation
The modified Cartan equation in compatible pairs:
@@
    This indicates that $\lambda$ is covariant constant relative to the connection $\omega$. For any horizontal vector field $X\in\Gamma(\mathcal{H})$:
@@
    Geometrically, this means that when translated along the horizontal direction of the connection, the change in $\lambda$ is exactly compensated by the adjoint action, ensuring that the constraint distribution $\mathcal{D}$ remains invariant under parallel transport. This generalizes the notion of covariant constant tensor fields to the context of constrained systems.
@@
    This ensures that compatible pairs $(\mathcal{D},\lambda)$ are preserved under gauge transformations, reflecting the mechanism of maintaining gauge symmetry in constrained physical systems.
@@
    Curvature Relations and Integrability Conditions: Let the curvature form be $\Omega=d\omega+\frac{1}{2}[\omega,\omega]$. Applying exterior derivative to the modified Cartan equation:
@@
    $0$ $=\mathrm{ad}^{*}_{\Omega}\lambda\quad(\text{when compatible pair exists})$
    This leads to the integrability condition $\mathrm{ad}^{*}_{\Omega}\lambda=0$, which is directly related to the Frobenius integrability of the constraint distribution $\mathcal{D}$. For any $X,Y\in\Gamma(\mathcal{D})$:
@@
    $=\langle\lambda,\Omega(X,Y)\rangle\quad(\text{since}\,\omega(X)=\omega(Y)=0\,\text{in}\,\mathcal{D})$
    $=0\quad(\text{from}\,\mathrm{ad}^{*}_{\Omega}\lambda=0)$
    Therefore $X,Y]\in\Gamma(\mathcal{D})$, showing that constraint distributions in compatible pairs are automatically integrable.
@@
    Momentum Mapping Interpretation: From a symplectic geometry perspective, $\lambda:P\to\mathfrak{g}^{*}$ in a compatible pair acts as a momentum mapping satisfying:
    $d\langle\lambda,\xi\rangle=\iota_{\xi_{P}}\omega_{\text{symplectic}},\quad\forall\xi\in\mathfrak{g}$
    where $\xi_{P}$ is the fundamental vector field generated by $\xi$. The modified Cartan equation corresponds to the equivariance condition of the momentum mapping, revealing the deep compatibility between $\lambda$ and the $G$-action on the principal bundle. In physical systems, this corresponds to the balance between constraint forces and gauge fields.
Remark 5 (Physical Significance).
The modified Cartan equation ensures that constraint distributions maintain structural consistency during system evolution, keeping constraint conditions dynamically coordinated with the geometric characteristics of the system. This mechanism explains why constrained systems satisfying the strong transversality condition can generate geometrically self-consistent dynamical behavior, especially in cases involving gauge fields.
3.2.2 Equivalence with Atiyah Exact Sequence Splitting